### v1

# Towards Foundation Models: Evaluation of Geoscience Artificial Intelligence with Uncertainty

Abstract
Artificial intelligence (AI) has transformed the geoscience community with deep learning models (DLMs) that are trained to complete specific tasks within workflows. This success has led to the development of geoscience foundation models (FMs), which promise to accomplish multiple tasks within a workflow or replace the workflow altogether. However, lack of robust evaluation frameworks, even for traditional DLMs, leaves the geoscience community ill prepared for the inevitable adoption of FMs. We address this gap by designing an evaluation framework that jointly incorporates three crucial aspects to current DLMs and future FMs: performance uncertainty, learning efficiency, and overlapping training-test data splits. To target the three aspects, we meticulously construct the training, validation, and test splits using clustering methods tailored to geoscience data and enact an expansive training design to segregate performance uncertainty arising from stochastic training processes and random data sampling. The framework’s ability to guard against misleading declarations of model superiority is demonstrated through evaluation of PhaseNet, a popular seismic phase picking DLM, under 3 training approaches. Furthermore, we show how the performance gains due to overlapping training-test data can lead to biased FM evaluation. Our framework helps practitioners choose the best model for their problem and set performance expectations by explicitly analyzing model performance at varying budgets of training data.
1 Introduction
As geoscience has breached the big data world within the last 20 years [3], the automation of geophysical tasks [^1] In the AI community, a task is a single defined goal that a model should accomplish, such as classifying an image or acting as a chatbot. through artificial intelligence (AI) via deep learning models (DLMs) is booming [44, 23, 27, 38, 26]. Foundation models (FMs), a subset of AI models trained on vast amounts of data in a self-supervised fashion that can be adapted for various downstream tasks using relatively smaller labeled datasets [4], are extending DLMs into new horizons and are actively being developed for geoscience [37]. The technological leap from traditional geoscience DLMs to FMs is inevitable, yet we still lack robust evaluation methods for even the simpler DLMs. As a result, we are unprepared to fairly compare and evaluate geoscience FMs. Without a robust evaluation approach, we are unable to determine whether new technological developments are truly state of the art, we lack diagnostics that assist developers in pinpointing areas for model improvement, and we fail to convey the scenarios under which a model can be trusted. In this research, we address this gap by identifying three fundamental evaluation aspects for AI that are directly applicable to current DLMs and extendable to future FMs, uniting them under a single evaluation framework.
The first and perhaps most important evaluation aspect is performance uncertainty. Two identical AI model architectures with identical training approaches will exhibit different performances as a result of different initializations and random batching during training [6, 35]. This is often referred to as aleatoric uncertainty. Here, we refer to it as training uncertainty or variability. Furthermore, the data sample and how data is split data into training, validation, and test partitions have drastic random effects on performance benchmarking [6]. This aspect is often referred to as epistemic uncertainty. Here, we refer to it as data uncertainty or variability. Determining algorithmic superiority in light of both sources of variation is a known issue [9, 16, 10, 40] and, to our knowledge, has not been explored for the geoscience AI community.
The second vital, yet often unexplored aspect helpful to researchers, and critical for FM comparison, is evaluating the model’s learning efficiency. Here, efficiency is defined as the relationship between the gain in performance to the additional training data required (see also [18]). Mapping this relationship helps decision makers compute the cost to benefit ratio of acquiring more data to meet a problem’s performance requirements, especially in data scarce scenarios [1]. It is particularly important to FMs because they are fine-tuned using smaller, targeted datasets for the task at hand, and the FM that best adapts may be preferred. This aspect may appear easy to measure by training two models on varying budgets of training data and comparing their performances. However, evaluating learning efficiency for geoscience domains requires thoughtful construction of these budgets, as simple random samples from geoscience datasets are often highly correlated due to geography, a limited number of measurement sites, and repeated measurements of a single source. This concept, known as data leakage, has been previously explored in the biological community [12, 33]. Removing or reducing data leakage to the extent possible when creating varying budgets of training data is crucial for unbiased estimation of learning efficiency.
The third fundamental aspect applicable to AI evaluation is the effect on model performance resulting from overlap in the training and test data. An overlap between training and test data is a clear example of data leakage that can lead to overestimation of model performance. To address this, it is common practice in the AI community to clearly partition the data into train and test sets such that evaluation on the test data reflects model’s performance on truly new (or unseen) data. However, ensuring no overlap between train and test data may be difficult for FMs, because a major component of the "foundation" in FMs is the big data used for pre-training that is often scraped from any available source. As an example, one of these sources could be the benchmark dataset that an evaluator is planning to use for testing. In comparing two FMs, one that happens to be trained on said benchmark and one that was not, we must avoid biased evaluation (see [11] for an example from the natural language processing community). Furthermore, spatial correlation in geoscience data warrants meticulous construction of training, validation, and test splits in contrast to simple random sampling for determining the effect in performance from overlap [17, 11, 12].
This research fundamentally contributes an evaluation framework for geoscience models that intertwines all three aspects: performance variation from data and training, learning efficiency, and train-test overlap that is directly applicable to current DLMs and intends to prepare the geoscience community for evaluation of FMs. We utilize concepts from statistical experimental design to measure each aspect in turn. We demonstrate our framework using a seismic DLM to directly compare 3 training approaches: training from scratch and two transfer learning approaches. We measure uncertainty through deep ensembles and random data selection processes, learning efficiency through careful data splitting and varying data amounts, and train-test overlap using overlapping and non-overlapping datasets. Finally, we cast all performance results into a statistical framework for direct comparison with uncertainty.
We use a conventional seismic DLM [^2] The evaluation framework was developed with specific challenges of FMs in mind. This development was forward-leaning and at the time of conducting this research, there are no geoscience FMS available for direct evaluation. Demonstration of the framework with a geoscience FM, particularly for evaluating downstream tasks, is a research path of future interest. for demonstrating the evaluation framework. A plethora of research in applying DLMs to the cardinal task of phase-picking provides a strong basis for our demonstration. Phase picking is the task of identifying and predicting the time that event signatures arrive, typically the primary and/or secondary phases (P-arrival and S-arrival, respectively), at a seismic station. Phase picking is a vital task for detecting, locating, and characterizing seismic events such as earthquakes or explosions. It can be a cumbersome task, motivating researchers to actively constructing better DL phase picking models (referred to as pickers here-on) [7, 28, 46, 5, 47, 48, 39, 31] and others are using these pickers to expedite workflows, including catalog creation [8, 32, 43, 2]. Of course, the best performing and most robust picker is preferred, but determining which picker is best for a given context remains a challenge that our framework is positioned to address.
The paper is formulated as follows. We first introduce the data sets and how the data splits are constructed to incorporate the three fundamental evaluation aspects. Then, we briefly introduce the model architecture and training scheme, including curating the inputs and targets, and provide a detailed review of the 3 training approaches and 5 models used for comparison. This is followed by the evaluation approach including the statistical framework that distills the different model performances for digestible comparison. Finally, we present the results in terms of performance and uncertainty and finish with concluding remarks and future work.
2 Data
Figure 1: An example waveform from the curated INSTANCE dataset showing the 3 component response with estimated P-arrival (labeled P-pick is the vertical blue line) and estimated S-arrival (labeled S-pick is the vertical red line) times.
We use two publicly available, labeled, non-continuous waveform datasets: the Stanford Earthquake Dataset (STEAD; [29]) and the Italian seismic dataset for machine learning (INSTANCE; [25]). STEAD is used for pre-training and INSTANCE for subsequent training and testing. STEAD contains $\sim 235k$ noise waveforms and $\sim 1.03M$ source waveforms representing $\sim 441k$ unique sources. For the INSTANCE dataset, we use the version measured in counts and only the earthquake waveforms measured in meters per second (mps). The resulting INSTANCE set includes $\sim 132k$ noise waveforms and $\sim 824k$ source waveforms arising from $\sim 54k$ unique sources. An example waveform from INSTANCE is shown in Figure 1 along with the labeled P-arrival and S-arrival. We access both datasets through the publicly available Python seismic deep learning package, SeisBench [42], which also provides a more detailed description of each dataset.
To measure the effect of overlap between training and test data on performance, we utilize the geographical overlap between STEAD (a global dataset) and INSTANCE (Italy only dataset). This results in a 3rd dataset which is a subset of STEAD we will refer to as STEAD Masked. STEAD Masked uses all of STEAD except the waveforms whose source or station locations overlap with INSTANCE’s area. We define INSTANCE’s area as the rectangle whose 4 corners are the pairings of the maximum and minimum source or station latitude/longitude with a 5 degree buffer in all directions. This rectangular mask reduces STEAD by 1.72% of the noise waveforms and 2.82% of the earthquake waveforms and reduces the number of unique sources by 0.97%.
2.1 Training, Validation, and Test Splits
Default splits provided in SeisBench for STEAD are used to pre-train our models. The STEAD masking process removes waveforms from each split at an approximately equal rate of 4%, so no bias is induced. A more meticulous process is implemented for INSTANCE data since it is used for both training and testing models. In the case of INSTANCE data, simply dividing waveforms or sources randomly into training, validation, and test partitions is inappropriate because data from nearby sources or stations share similar sources, geographies, and processes, and hence, are correlated. Therefore, if we randomly sample the data on a waveform or source basis, we inadvertently add information about the other waveforms or sources within the dataset that were not selected. This results in data leakage, a challenge described in the introduction that impacts evaluation of learning efficiency as well as training and test overlap. To address this, we divide the INSTANCE dataset into spatial clusters, a simple and strong indicator of signal commonality. We have the option to cluster earthquake waveforms based on source or station locations and we choose to cluster based on source locations for three reasons:
  1. 1.
    sources are most often detected by sensor stations nearest to the source,
  2. 2.
    waveforms sharing the same source are assumed to have more in common than waveforms sharing the same sensor,
  3. 3.
    and the sources tend to group themselves based on geographic contours across Italy, thereby making a clustering approach natural.
We use a k-means algorithm from the Python package, Scikit-learn [34], to cluster earthquake sources based on their latitude and longitude. We chose $k=20$ clusters because it lead to clusters large enough to feasibly represent the ground characteristics of their geographic region, yet small enough to explore model performance in limited data settings. To assign the noise waveforms, which have no corresponding source locations, to the established clusters, we use the trained k-means model to predict its associated cluster based on its station location. The clusters and number of sources, stations and waveforms are shown in Figure 2. Table 1 summarizes the amount of data in the training, validation, and test splits and the construction of the splits are described next.
Figure 2: INSTANCE data with 20 colored clusters for the sources (top, circles) and the stations (bottom, triangles). For the sources, the numbers indicate the number of sources in thousands followed by the number of waveforms in thousands. For the stations, the numbers show the number of stations and the number of noise waveforms in thousands. The colored text outline designates the group: north test group is cyan, training/validation group is white, and south test group is red.
To construct the INSTANCE splits, we first segregate the clusters into two groups: a training/validation pool and a test pool. We select the 12 central clusters to be the pool from which the training/validation data will be constructed, and 8 clusters - the 4 northern most and 4 southern most - to be the test clusters (see Figure 2). Choosing the test clusters from northern and southern regions allows us to explore two different concepts. First, they help to evaluate model performance on semi-out-of-distribution data: we expect the waveforms to share similar signatures as a whole because they are regionally constrained, but they should be slightly different due to more localized geological distinctions between the areas. Second, these splits represent a realistic situation where we may know how a model performs in a certain region and we want to measure its consistency for neighboring regions.
To mitigate regional bias in the the INSTANCE test split, we balance the data from north and south regions. The northern region is smaller comprising 2,269 sources, and therefore we randomly sample 2,269 sources from the southern region. For each source selected in the test split, we include all associated waveforms, resulting in a test set totaling 49,516 earthquake waveforms: 23,086 from the north and 26,430 from the south. To maintain the overall ratio of 11.4% noise to earthquake waveform ratio in INSTANCE, we randomly select 2,822 noise waveforms from each of the south and north regions.
Table 1: The data splits are summarized showing the number of sources, earthquake (EQ) waveforms (WFs), and noise WFs. For the training split, values should by multiplied by the quantity of clusters used in the training design.
Data Split Sources EQ WFs Noise WFs
Training 636 $\times$ # of clusters 8,878 $\times$ # of clusters 1,106 $\times$ # of clusters
Validation 1,908 27,169 3,312
Test 4,538 49,516 5,644
We aim for an 80-20 ratio for the training to validation splits and we desire an equal number of sources from each cluster to reduce bias. Since the smallest of the 12 central clusters has 795 sources, we randomly sample 159 sources from each cluster for the validation set. This totals to 1,908 sources encompassing 27,169 earthquake waveforms. Maintaining a $\sim 11.4\%$ earthquake to noise ratio, we then randomly sample 276 noise waveforms per cluster totaling 3312 noise waveforms for the validation split.
After removing the INSTANCE validation set from the central clusters, the remaining waveforms become the pool from which we build the training splits. When training models, the amount of training data is specified by the number of clusters used for training. However, to keep data amounts relatively constant across clusters, each time a cluster is selected to be included in training data, we randomly select 636 sources (80% of 795, the smallest number of sources in any cluster) which on average results in 8,878 earthquake waveforms, and 1,106 noise waveforms per cluster.
3 Methods - Model Architecture, 3 Training Approaches, and 5 Models
Table 2: The details of the 5 models are given in each row including their names, if the model is initialized randomly or using weights from an upstream model trained on either STEAD or STEAD Masked, and whether weights were frozen.
Name Initialization Weights
Standard Random Free
TL Free STEAD Free
TL Free Masked STEAD Masked Free
TL Frozen STEAD Half Frozen
TL Frozen Masked STEAD Masked Half Frozen
We develop a unifying framework for evaluating AI models that considers uncertainty, learning efficiency, and training-test overlap to guide model selection. We demonstrate our framework by comparing 5 models as outlined in Table 2 to compare the effects of 3 model training approaches: training from scratch (the standard DL approach) and two transfer learning (TL) approaches with varying number of free parameters. However, the framework presented is sufficiently broad where one could explore the effects of different hyperparameters, training approaches, or new architectures altogether. We first review the model architecture and other commonalities among the models, then we review the uniqueness of each model and how the training design allows for estimation of data and training uncertainty.
All models use the PhaseNet [47] deep learning architecture which has been extensively utilized for phase picking [20, 5, 30] and features a U-Net type architecture [36]. PhaseNet intakes 3 component waveforms of length 3,001 and is trained to output probabilities of P-arrivals, S-arrivals, and noise for each sample (time point in the waveform). Since our goal is to demonstrate the evaluation framework, we focus only on the P-arrival predictions and discard the S-arrival and noise predictions for all subsequent analysis.
All models use the PhaseNet [47] deep learning architecture which has been extensively utilized for phase picking [20, 5, 30] and features a U-Net type architecture [36]. PhaseNet intakes 3 component waveforms of length 3,001 and is trained to output probabilities of P-arrivals, S-arrivals, and noise for each sample (time point in the waveform). Since our goal is to demonstrate the evaluation framework, we focus only on the P-arrival predictions and discard the S-arrival and noise predictions for all subsequent analysis.
To demonstrate our evaluation framework, we consider 3 training approaches: the standard approach for deep learning and two transfer learning approaches. The standard approach entails randomly initializing model weights and pulling the training, validation, and test splits all from the dataset of interest, in this case INSTANCE. Thus, the standard model represents this approach.
The two transfer learning approaches are thought to be useful when a small amount of labeled data is available from the region of interest, but a large amount of labeled training data can be obtained from other regions. As such, one first trains a model with data from the other regions (we call an upstream model), then switches to data from the region of interest to continue learning, thereby transferring the accumulated knowledge to the new domain. In our case, we first train PhaseNet using STEAD or STEAD Masked and then further train using INSTANCE. We explore two options: 1) allow all weights to change during INSTANCE training phase and 2) fix half of the weights (the input and down sampling layers). The TL models that have fully free weights are denoted as Free, and those with fixed weights are denoted as Frozen.
To explore the effect of training-test overlap of interest to FM evaluation, we train two models for each of the two TL approaches by changing whether the model was initialized from the model trained on STEAD or STEAD Masked. Those that are seeded by STEAD Masked have the Masked label. This results in 4 TL models named TL Free, TL Free Masked, TL Frozen, and TL Frozen Masked. All 5 models (the Standard model and 4 TL models) are validated and tested using INSTANCE data.
To evaluate learning efficiency, we train each model using 1, 3, 6, 9, and 12 clusters from the INSTANCE training clusters. To measure data uncertainty at a given quantity of clusters, we randomly select the cluster set 12 times, training the model each time. Then, to separate the data uncertainty from training uncertainty, we use deep ensembles [21, 14, 19]. That is, for each of the 12 cluster sets, we retrain the model 4 different times under different initializations (holding the training data sample fixed) to obtain 4 model instances for evaluation. To obtain different initializations for the Standard model, we simply randomly reset the weights. For the TL models which are seeded by weights from an upstream model, we first train 48 upstream models for both STEAD and STEAD Masked starting from random initializations. Then, we initialize the TL models by pulling the set of weights from one randomly selected upstream model.
We use ensembles to obtain training uncertainty for a few reasons. Fundamentally, deep ensembles are easy to understand and implement and they only require extending training time. This is in contrast to other uncertainty methods which generally attempt to approximate the posterior distribution of the prediction. These methods include Bayesian neural networks [22] and approximations thereof including dropout [13], Stochastic Weight Averaging - Gaussian (SWAG) [24] and many others [15]. However, these approaches enforce distributional assumptions of the model weights, make strong/arbitrary decisions (e.g, dropout), rarely consider different basins of attraction, and are generally more difficult to interpret. Some methods, however, such as MultiSWAG [41] (see [2] for a seismic based implementation), offer a strong balance between the complex Bayesian assumptions and the simpler empirical ensembles and merit consideration for future work.
In our deep ensemble implementation, we loosely follow the snapshot ensemble approach of [19] which boasts computational advantages in finding different parameter states that minimize loss. We implement a cosine annealing learning schedule with 4 restarts with an initial learning rate of $0.01$ and final rate of 00. These learning parameters exhibited the best performance in preliminary training experiments especially for models with the smallest quantities of clusters (one cluster). However, simple rate resets routinely failed to push models with small training sets into new basins of attraction, so we altered the approach towards the classical deep ensemble method [21] by fully reprogramming the model weights at each scheduled rate reset rather than relying on rate resets alone. Depending on the model being trained, this reprogramming may be a complete randomized reset (for the OOD and Standard models) or replacing weights by randomly pulling from the 48 upstream STEAD (Masked) models (for the TL models).
For ease of discussion, we define a single "model instance" as the combination of model, quantity of clusters, cluster set (random selection of those clusters), and initialization (ensemble). For example, a single model instance of the Standard model could be trained using 3 clusters randomly chosen - in reference to Figure 2 - to be clusters red, blue and light-green, and the third initialization. With 4 TL models and 1 Standard model each trained using 5 quantities of clusters over 12 random cluster sets with 4 initializations, we obtain 1,200 model instances.
We train all models using a sufficient number of epochs until convergence. In general, the Standard models reached convergence within 50 epochs and the TL models reached convergence within 12 epochs. We selected the model instance for evaluation that exhibited lowest loss on the validation data within a cosine annealing cycle.
4 Evaluation
4.1 Metrics
As our main goal is to demonstrate the effectiveness and versatility of our evaluation framework, we limit the scope of evaluation of the phase picking models by considering just a few classification metrics and one functional based regression metric. To form these metrics, we first implement techniques similar to those found in [47, 30, 20] to obtain model predictions for test waveforms. The predictions are probabilities of the P-arrival for each sample in each waveform. From the predicted curves and through an algorithm similar to those found in [ref same papers as above], we obtain P-arrival picks, which are the discrete locations in time that a P-arrival is predicted by the model. There can be none, one, or any number of picks per waveform. To motivate the specific metrics we consider, we note that it is important in evaluation of phase pickers to consider both the model’s strength in identifying if a P-arrival occurs (e.g., alerting) and, if so, the associated accuracy of the pick (e.g., how close is the predicted pick to the true P-arrival).
Evaluating identification is typically done using classification metrics defined using the picks [47, 30, 20, 45]. True positives ($TP$) are assigned to picks if they are near enough to the labeled P-arrival. False positives ($FP$) are assigned to picks either far from the labeled P-arrival or to extra picks that are not true positives. False negatives ($FN$) occur if no picks occur near the labeled P-arrival. True negatives are not informative in this context and are not recorded. In our display, we consider summary metrics including $Recall=\frac{TP}{TP+FN}$ and $F1=\frac{2TP}{2TP+FP+FN}$ which are only applied to earthquake waveforms. An example of a model prediction featuring classified picks including a true positive and 3 false positives is shown in Figure 3. For noise waveforms, we calculate noise percent correct as the number of noise waveforms containing no predicted picks ($TN_{noise}$) divided by the total number of noise waveforms ($n_{noise}$). That is, $Noise\,\%\,Correct=\frac{TN_{noise}}{n_{noise}}$.
Evaluating accuracy is typically dependent on the classification results and done using regression-based metrics [47, 30, 20, 45]. If a pick is classified as a true positive, then we obtain a residual defined as the difference in time between the labeled P-pick and the predicted P-pick. Because different models can produce different numbers of true positives, typical regression metrics like the root mean square residual (RMSR; the square root of sum of the residuals divided by the number of true positives) fail to convey the full story. Therefore, functional based metrics like cumulative RMSR, defined as RMSR calculated using only the residuals whose magnitudes are below some value, are more informative.
Figure 3: The predicted probability of the P-arrival for the waveform shown in Figure 1. The labeled P-arrival is displayed on the x-axis as a blue triangle along with 4 predicted picks (green stars) along with labels that classify the pick as either a false positive or true positive. An inset plot zooms in on the labeled pick to show the true positive and an extra nearby pick classified as a false positive.
4.2 Statistical Framework for Evaluation
In this section, we present a statistical framework for comparing models in terms of their performance and uncertainty arising from data sampling and model training. The statistical framework can be thought of as a mixed effect model or a completely randomized design with sub-sampling. It allows us to distinguish performance effects with uncertainty between the 5 models and different training amounts. Furthermore, we can segregate and estimate the two sources of variation with uncertainty. We describe our overall approach here and relegate many details of the calculations to the appendix.
For a given metric that summarizes the outcomes of a model instance, denoted $y_{madi}$, where $y$ is the metric (e.g., Recall), we setup the statistical model
$y_{madi}=\gamma+\mu_{m}+\alpha_{a}+\theta_{ma}+\epsilon^{data}_{mad}+\epsilon^{train}_{madi},$ (1)
where $m=\{1\ldots 5\}$ indexes the 5 models, $a=\{1,\ldots,5\}$ indexes the 5 quantities of clusters (representing the amount of data) at levels $\{1,3,6,9,12\}$, $d=\{1,\ldots,12\}$ indexes the randomly selected cluster set (i.e., the data choice selected to train the model), and $i=\{1,2,3,4\}$ indexes the 4 model initializations.
In this statistical model, $\gamma$ is the grand mean, $\mu_{m}$ is the effect of model $m$, $\alpha_{a}$ is the effect of the $a$’th quantity of clusters, and $\theta_{ma}$ is the interaction term. The uncertainty associated with data selection and model training is segregated into two independent random terms each assuming a zero mean independent and identically distributed Gaussian with variance inherent to a given model and quantity of clusters: $\epsilon^{data}_{mad}\sim\mathcal{N}(0,\sigma^{2data}_{ma})$ and $\epsilon^{train}_{madi}\sim\mathcal{N}(0,\sigma^{2train}_{ma})$. This is reasonable because we desire to understand how the uncertainty is impacted by the model choice and quantities of clusters used.
Using these models, we estimate the effects on the metric for each model and quantity of clusters by computing the mean over the 12 cluster sets and the 4 initializations, denoted as $\bar{y}_{ma\cdot\cdot}$. Then, we compute $90\%$ confidence intervals around those estimates using the Gaussian assumptions. Note that the intervals are influenced more heavily by the cluster set (data choice) uncertainty because at each cluster set, we average over the initializations thereby reducing training uncertainty influence.
The above statistical model and resulting confidence intervals rely on Gaussian assumptions for the two error terms. That is, the distribution of the error terms implies a distribution of the metric as $y_{madi}\sim\mathcal{N}(\mu_{m}+\alpha_{a}+\theta_{ma},\sigma^{2data}_{ma}+\sigma^{2train}_{ma})$. We recognize that metrics such as recall are bounded between $[0,1]$, so the Gaussian assumption theoretically cannot hold. However, in practice we have a large number of test waveforms to form these metrics and we find that the variances of these metrics are small compared to the range of the metrics’ bounds.
5 Results
After training the 1,200 model instances, computing the metrics for evaluation, and organizing them through the statistical framework, we are ready to compare the performance of the models.
The classification based summary metrics of recall and noise percent correct are presented in Figure 4 along with 90% confidence intervals for the standard and 2 masked TL models showing their performance as a function of quantity of clusters used in training. All models improve in recall as more data is added with largest gains between 1 and 3 clusters followed by diminishing returns from there on. The trend is less consistent for noise percent correct where the TL w/ Frozen Weights Masked model performs slightly worse with increasing data. The variability in performances decreases for all models as more training data is added.
Figure 4: The statistical framework is used to summarize classification results including earthquake recall and noise percent correct as a function of the number of training clusters. Vertical bars represent 90% confidence intervals. Largest performance gains occur as the quantity of clusters increases from 1 to 3 and crossing lines indicate some training approaches are superior for different budgets of training data.
The functional regression based metric of cumulative RMSR demonstrating the evaluation’s framework versatility and importance is shown in Figure 5 for the standard and 2 TL masked models with 90% confidence intervals (ribbons) for selected quantities of clusters of 1, 6, and 12. Stark performance differences can be seen between each model especially as more training data is added as seen through tightening confidence intervals.
In both Figures 4 and 5, we see that there is no clear global winner in terms of learning efficiency. The TL w/ Free Weights Masked model appears superior for recall and cumulative RMSR for the smallest training data size of 1 cluster, but is surpassed by the standard model for greater than 3 clusters. In contrast to expectations, adding more data actually degrades the TL w/ Frozen Weights Masked model for noise percent accuracy. These results would help practitioners choose the model for their given problem. Depending on the availability of training data and how they prioritize performance metrics, their choice would change.
Figure 5: The statistical framework is used to summarize the regression-based functional metrics of cumulative RMSR for 1, 6, and 12 clusters (top, middle, and bottom respectively) for the standard and 2 TL masked models. Ribbons indicate 90% confidence intervals. The relative positions of the model curves change as more training data is added indicating that no model is a globally most efficient learner.
Valuable to AI evaluation is determining whether there is an effect of overlap between training and test data. To explore this aspect, we compute statistical contrasts (differences) between the unmasked and masked models. We show these contrasts in Figure 6 computed for recall, F1, and noise percent correct for the two TL models at each data amount with 90% confidence intervals. Vertical lines are drawn at the origin and if the interval contains 00, then the effect of overlap may be considered statistically significant at the 10% level. Positive values indicate that including overlap during training improves model performance.
In analyzing Figure 6, we see minimal impact from overlap for 1 cluster, but loosely positive trends with increasing training data for recall and F1. For noise percent correct, however, there appears to be little impact of overlap regardless of the training data amount. These results demonstrate that overlap should be considered in future DLM/FM evaluation because overlapping may inflate performance metrics and fair evaluation should account for such bias.
Figure 6: Contrasts between Unmasked and Masked models with 90% confidence intervals shown for the TL w/ Free Weights (red-orange; top 3 panels) and TL w/ Frozen Weights (sky-blue; bottom 3 panels) models for recall (left), F1 (center), and noise percent correct (right) evaluated at each quantity of cluster. The figure shows modest performance gains as a result of training-test overlap.
Our evaluation framework allows for estimation of two sources of performance variation including random data sampling and stochastic training processes. As an example, Figure 7 shows these estimates for F1 for the standard and 2 TL masked models across the quantities of training clusters with 90% confidence intervals. We see that training and data variance decrease as more training data is added and we see some differentiation in the variability for each model. The data variance (middle panel) shrinks quickly towards zero as more data is added and is often estimated as negative. These negative estimates are not shown because negative variance estimates are theoretically impossible. This is a known challenge in these classical statistical frameworks, but negative values essentially mean the data variance component is very small relative to the training variance.
The final panel in Figure 7 shows the relative contributions of each source of variation through bar plots where the opaque and transparent bars are the contributions of data and training variances, respectively. The total variance for all models trends downward with more data with diminishing returns. The variation is mostly dominated by the stochastic training processes for quantities of clusters greater than 3. Practitioners can use this information to target specific sources of variation that they suspect might dominate in their problem, perhaps by running more initializations or labeling data more strategically.
Figure 7: Training variance (left) and data sampling variance (middle) are estimated with 90% confidence intervals (vertical bars) for F1 for the standard and 2 TL masked models. The right panel shows bar plots constructed as the sum of the estimates of training and data variances. The different transparency levels show the contribution of each component where the training contribution is transparent and the data contribution is opaque. Estimating performance variation and attributing the source is informative to practitioners during model selection for setting performance expectations and risk mitigation.
We provide an example that emphasizes the risk of evaluating models without considering variation in performance. We consider the scenario where a practitioner is deciding among 3 models (the standard model and the 2 TL masked models) to incorporate into their workflow. They are equipped with a single cluster set, say 1 cluster’s worth of data. To make their decision, we imagine that they train one instance of each model using the fixed data. We did exactly this in our training design, but we went further by training each model 4 times under different initializations, not just once. As such, we imagine the practitioner is equally likely to get one of the four recall scores we obtained for each model. Thus, with three models and 4 equally likely results for each, there are $644^{3}=$ 64 equally likely possible outcomes that the practitioner could see. If the practitioner ranks each model based on recall score (e.g., first place, second place, third place), due to random training variation, we might expect these rankings to change if the practitioner ran the experiment a second time. The important question then becomes, what is the probability that each model gets each rank?
Using our 3 models each initialized 4 times for a given quantity of clusters, we calculate the exact probabilities that each model is ranked in first, second, or third place. Furthermore, since we have 12 cluster sets for each quantity of clusters, we conduct the calculation $12$ times and take the mean to mitigate data bias. The results of this simulation are presented in Figure 8, where the probability estimates of ranking each model based on recall in first, second, or third place are shown for $1$ and $12$ cluster quantities.
Figure 8 shows that, due to training and data uncertainty, there is substantial risk of choosing a model that performs more poorly on average. For example, for 1 cluster (top panel), there is a $1.0\%$ chance that the TL w/ Frozen Weights Masked (orange) model has the best recall score. In light of the recall results in Figure 4 (left panel) for 1 cluster, we know that the TL w/ Frozen Weights Masked performs substantially worse than the other two models on average. Based purely on recall, it would be a mistake if the practitioner were to choose this model for their workflow. In a similar light, claiming that the TL w/ Frozen Weights Masked model is the state of the art compared to the other two would be an even graver mistake. Our framework guards against such mistakes.
We show the probability of rankings results for the 1 and 12 clusters to show how the probability of mis-ranking is a greater risk in the lower data regime and the risk is somewhat reduced as more data is added. From these results, we learn that researchers should apply the same principles in our framework when evaluating FMs. This is because FMs are expected to need minimal labeled training data, but teasing out exactly how much data is needed to avoid the risk of mis-ranking aids in determining which model is best.
Figure 8: To demonstrate the importance of considering uncertainty in the evaluation framework, the probability of ranking the standard and 2 TL masked models based on recall in first, second, or third place is shown for 1 (top) and 12 (bottom) quantities of clusters. Numbers/bars indicate the probability of ranking the associated model in each place if one were to only train one model instance. Numbers less than 100% indicate that mis-ranking the models is a risk which is detrimental to model selection.
6 Discussion
We present results that will be increasingly relevant for developers of DL pickers and for the seismic community as they invest in FMs. In inspecting the results presented in Figure 4, we see overall poor performance in the TL w/ Frozen Weight models. We hypothesized that freezing half the model at the learned weights from the global dataset would produce the best model for small training data amounts. However, the results negate this hypothesis for datasets comprising 10,000 waveforms (approximately the number of waveforms in a single cluster) or more for the PhaseNet architecture. In a similar vein, we see that models trained from scratch perform better than TL when the data size approaches 30,000 waveforms (3 clusters). Since FMs are pre-trained using a large general dataset and fine-tuned (typically weights for the last couple of layers are set to be updated) using tailored datasets to accomplish downstream tasks, we see the importance of continuing to assess the utility of seismic FMs by comparing to traditional modeling approaches.
Another valuable result for the seismic AI community is that expending effort to acquire larger and larger labeled training datasets may be wasteful. This is exhibited by Figure 4 where we see diminishing improvement as the amount of training data increases. This suggest relatively smaller datasets ($\sim$30,000 or $\sim$60,000 waveforms) may be sufficient for reasonable model performance even when training models from scratch. This is a good news because large labeled datasets are expensive and sometimes unattainable. Furthermore, these results fortify the importance of learning efficiency as an aspect of future FM evaluation. Since the best model depends on the amount of training data, knowing which model or modeling approach does more with less is valuable information during model selection. And, as we have shown, it is not a simple one-model-takes-all, but rather a dynamic choice for a given situation.
As we work toward an evaluation framework with uncertainty for geoscience FMs, we recognize the limitations in using deep ensembles as the primary form for obtaining uncertainty estimates. Deep ensembles are computationally expensive, requiring at least some retraining. With FMs, unfortunately this expense is likely prohibitively large for retraining the base model. However, this limitation does not detract from the importance of quantifying training induced variation, and alternative approaches are possible such as using ensembles during fine tuning.
We have demonstrated evidence that overlap in the training/test data can imbue performance gain, but we recognize that the effect may be confounded by the 4% decrease in the amount of data in the STEAD dataset with Italy removed. In other words, it is unclear if the performance drop in the masked models are caused by removing data from the Italian region, or if the drop is simply caused by the withholding of data from any region. To tease out these confounded effects, we could train models using random regional masking, then compare their performances to the performance of the Italy masked models. If the Italy masked models are significantly worse, then we may more confidently conclude whether the cause is the Italian region.
Regarding constructing splits to measure learning efficiency, location based clustering of the INSTANCE data was a convenient and effective first order attempt for making clusters that represent different aspects of the INSTANCE data that can be used for future frameworks. However, other methods that can characterize and group the INSTANCE data based on some inherent characteristics/features could better limit leakage between the clusters. These methods could utilize other metadata such as source to station distance, earthquake type, geological characteristics of the area, or signal-based aspects. A potential drawback of these methods is that they may result in fewer clusters, thereby reducing the number of cluster sets and making the data variance more difficult to estimate.
Related to data variance and clustering is that some data variances are estimated as negative because they are small relative to the training variance. One way to more precisely estimate data variance is to increase the number of cluster sets (i.e. from 12 randomly chosen cluster sets to 30). However, we would not see much variation in the cluster sets for large quantities of clusters (e.g., 9 or 12) because we do not sample with replacement. Sampling with replacement would allow larger cluster quantities to have more variety in their cluster compositions. But, this also has a drawback in that some models may get biased towards a specific cluster if it appears multiple times in the cluster set and may not best represent data sampling variability. Another option is to create more clusters such that each cluster set is more likely to be a unique combination of clusters and the probability of a single cluster appearing multiple times is diminished. However, we are limited in decreasing the cluster sizes because as the total number of clusters approaches the number of sources, the issue of data leakage resurfaces. Determining the best solution to this issue is a fruitful topic for future research.
There are also avenues for improvement for the statistical framework. Since many of the models share similar characteristics, for example Free or Frozen weights, masking or no masking, etc., we could improve at consolidating this information in the model. Bayesian hierarchical modeling, while more complicated, would allow for flexibility of this nature and could simultaneously address the negative data variance issue through priors.
7 Conclusion
We present an evaluation framework for current DLMs and future FMs that simultaneously considers performance uncertainty/variability, learning efficiency, and overlap between training and test data. We demonstrate our evaluation framework using PhaseNet, a popular phase picking DLM [47] to compare 3 training approaches (training from scratch and two transfer learning approaches). We implement an extensive training design comprising 5 models and 1,200 model instances over varying amounts of data from meticulously constructed data splits. We cast the results into a statistical framework which allows for direct comparison among models and training approaches across the training budgets.
We demonstrate the effectiveness and versatility of the evaluation framework by comparing classification and functional regression metrics. We show how no model is a global winner in terms of learning efficiency; that it depends on the total amount of training data used. We further demonstrate that training-test overlap must be considered in future FM evaluation as seen through performance improvement due to overlap. By providing estimates for the data and training uncertainty, a developer or practitioner can better target their efforts to mitigate sources they deem most important for their problem. We show that without the evaluation framework, there is a significant risk of model mis-ranking. Finally, our results demonstrate how transfer learning may not always be the best choice in limited data scenarios, which emphasizes the importance of continuing to compare FMs to their simpler DLM counterparts and inspect how training data budgets influence model performance. We hope that this research helps practitioners and developers by providing a general guide for what a truly revolutionary model might look like.
8 Acknowledgements
This work was performed under the US Department of Energy, National Nuclear Security Administration’s Office of Defense Nuclear Nonproliferation Research and Development. This work is approved for public release under LA-UR-24-31133. We would like to thank Yongsoo Park (Los Alamos National Laboratory), Lisa Linville (Sandia National Laboratory), and Leanna House (Virginia Tech) for their advice and/or edits.